OpenStudy (aravindg):

9. Two towns A and B are connected by a regular bus service with a bus leaving in either direction every T minutes. A man cycling with a speed of 20 km/h in the direction A to B notices that a bus goes past him every 18 min in the direction of his motion and every 6 min. in the opposite direction. What is the time period T of the bus service and with what speed(assumed constant) do the buses ply on the road?

OpenStudy (aravindg):

cnfused

OpenStudy (jamesj):

Let the bus travel at a speed s. Then the distance between the buses going from A to B is sT Now, if the man sees them every 18 minutes, how far has he travelled in 18 minutes?

OpenStudy (aravindg):

st/something

OpenStudy (aravindg):

hey wait

OpenStudy (jamesj):

He has travelled speed . time = 20 km/hr . 18/60 hr = 20 . 6/20 km = 6 km. Now for buses going from A to B, what is their velocity relative to the man?

OpenStudy (aravindg):

oh i got that part

OpenStudy (aravindg):

:)

OpenStudy (aravindg):

20-s

OpenStudy (aravindg):

james its 1:30 am here i am sleepy

OpenStudy (jamesj):

It is s - 20. Given that, find an equation that relates the speed of the bus s and the time T. Now find another equation for the buses traveling from B to A. If you do, you'll have equations in two unknowns---s and T---which you can now solve. === Come back in your morning then.

OpenStudy (aravindg):

k thanx

OpenStudy (jamesj):

I'll give you the first one. In the 18 minutes = 3/10 hour, the bus travels a distance of 3s/10. Now at the moment that one bus passes the man, the next bus has to travel a distance of sT + 6 to catch up. Draw yourself a diagram and convince yourself of all that. In which case: sT + 6 = (3/10)s